Waves of same amplitude and same frequency from two coherent sources overlap at a point. The ratio of the resultant intensity when they arrive in phase to that when they arrive with 90degrees phase difference is?

Grade:12

1 Answers

Eshan
5 years ago
Let the intensity of each wave beI.

Total intensity when two waves interfere with some phase angle\phiis

I_{net}=I_1+I_2+2\sqrt{I_1}\sqrt{I_2}cos\phi
Hence ratio of intensities in the two cases=

\dfrac{I+I+2Icos0^{\circ}}{I+I+2Icos90^{\circ}}
=\dfrac{4I}{2I}=2